OpenStudy (anonymous):

If the diameter of a cylinder is 14 feet and its height is 4 feet, the volume of the cylinder is 2,461.76 square feet. True False

OpenStudy (anonymous):

Diameter of a cylinder is 14ft Radius of cylinder is 14/2 = 7ft. Volume of cylinder = area of circle * height = pi * r^2 * h =pi * 7^2 *4 =615.75 ft ^3 False. (volume cant be in square feet anyway.

OpenStudy (anonymous):

thanks!